9. The Congressional Research Service provides information on the composition of Congress. For the 114th Congress. 18.7% of the members are senators (I hope this doesnt change too quickly!) and 46% of the senators are Democrats. What is the probability that a randomly selected member of the 114th Congress is a Democratic senator? 10. The National Center for Education Statistics publishes data about educational institutions in the Digest of Education Statistics. According to that publication, 13.4% of all students attend private schools, and 4.5 % of all students attend private colleges. What percentage of students attending private schools are in college? thank you!
Answer #1

9)P(democratic senator)=P(senator)*P(democrat|senator)=0.187*0.46=0.08602

10)

\P(% of student attending private schools are in college)=(0.134*0.045)*100=0.603%